OpenStudy (tetyana):

If we drop a small metal sphere into the water and watch its motion, it will quickly reach terminal velocity. If that final velocity is 5 cm/s, the radius of the sphere is 5 cm, the density of water is 1000 kg/m3, and the viscosity of the water is 10-3 Ns/m2, what is the Reynold's number of this sphere?

OpenStudy (mos1635):

I suppose you have a relative formula....

OpenStudy (tetyana):

What do you mean?

OpenStudy (mos1635):

a formula for the force a liquid aplieys to object mooving in ti it....?

OpenStudy (anonymous):

use the following formulae: \[N=(\rho \times v \times D )\div \eta\] where \[\rho\] is density \[v\] is terminal velocity.\[D\] is diameter.\[\eta \] is the viscocity of the liquid and N is reynold's number . find N.

OpenStudy (mos1635):

@taufique i feel there something wrong with your initial formula....

OpenStudy (mos1635):

N should be propotional to η

OpenStudy (anonymous):

@mos1635 no N will be inversely proportional to

OpenStudy (anonymous):

eta.

OpenStudy (anonymous):

Viscous Force F is given by \[F = 6\pi \eta Rv\] \[\eta\] = coefficient of viscosity R = radius of sphere v = terminal velocity From this we get \[\eta = F/6\pi Rv\]

OpenStudy (ash2326):

@suyash011 would you explain more?

OpenStudy (mos1635):

η=F/6πRv some thing like that i had in mind too..... but keep in mind that i have no idea who Reynold is, or his number...:) :) :)